OpenStudy (anonymous):

want easy method 4 this ques; rs 430 is divided among a,b,c in such a way thst if a gets rs2 b gets rs 3, & b gets rs 5,c gets rs 6. find b share ?

OpenStudy (anonymous):

let a gets 2x so b=3x,c=3x*6/5 so 2x+3x+18x/5=430 43x/5=430 x=50 b gets 3*50=150

OpenStudy (anonymous):

how did u get c=3x*6/5 ??

OpenStudy (anonymous):

dipan ???

OpenStudy (anonymous):

because if b gets 5 then c gets 6 so if b gets 1 then c gets 6/5 if b gets 3x then c gets 6*3x/5=18x/5 this is an unitary method.......

OpenStudy (anonymous):

okkk !! thnxx !! may i noe wat r u upto in ur life (wat u do, nd ur details ) ?

OpenStudy (anonymous):

what do u do...tell me first...

OpenStudy (anonymous):

okk., i am from india, nd i have completed my intermediate, prparing for my B.com honours entrance exam !

OpenStudy (anonymous):

i'm also from india...

OpenStudy (anonymous):

m an engineering student...

OpenStudy (anonymous):

form india ?

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

from*

OpenStudy (anonymous):

oh great ! which state ?

OpenStudy (anonymous):

bengal...nd u?

OpenStudy (anonymous):

hyderabad !

OpenStudy (anonymous):

good...

OpenStudy (anonymous):

how r u so perfect ??

OpenStudy (anonymous):

??

OpenStudy (anonymous):

why ??? am i perfect... in what sence...

OpenStudy (anonymous):

ya re ! u answering such quest !! so sayin u r perfect nd am thinkin y cant i be like u (in knowledge wise ) :p

OpenStudy (anonymous):

u will be....

OpenStudy (anonymous):

but any reasons for such intelligent metallity ?? :D

OpenStudy (anonymous):

i can't solve accountancy problem..that u can